### v5

# The conformal Laplacian and positive scalar curvature metrics on manifolds with boundary

Abstract.
We give examples of spin $4$-manifolds with boundary that do not admit metrics of positive scalar curvature and minimal or even mean convex boundary. These manifolds in fact have the stronger property that the conformal Laplacian with appropriate boundary conditions is never positive. The obstruction to the positivity of the conformal Laplacian is given by a real-valued $\xi$-invariant associated to the APS theorem for the twisted Dirac operator. We use analytic techniques related to the prescribed scalar curvature problem in conformal geometry to move beyond earlier work where the metric is a product near the boundary.
The second author was partially supported by NSF Grant DMS-1928930 while he was in residence at the Simons Laufer Mathematical Sciences Institute (formerly known as MSRI), as well as NSF FRG Grant DMS-1952790.
Math. Subj. Class. 2020: 53C21 (primary), 35J66, 53C27, 58J05, 58J20 (secondary).
1. Introduction
There is a well developed theory of determining which closed manifolds $M$ have topological obstructions to admitting positive scalar curvature (psc) metrics (see [27]). The corresponding theory for manifolds with boundary $(M,\partial M)$ is less complete. The main results in this paper are both geometric and analytic extensions of known results for psc metrics which are a product near the boundary. On the geometric side, we weaken the product metric assumption to the totally umbilic boundary case. On the analytic side, we prove results relating psc metrics to the positivity of the conformal Laplacian with certain geometric boundary conditions. The geometric and analytic results are related: for example, an obstruction to a metric with positive conformal Laplacian and minimal boundary is automatically an obstruction to a psc metric with the same boundary condition. As a result, the main technical work in the paper is in geometric analysis, specifically in conformal geometry on manifolds with boundary.
To state the main results, let $R_{g}$ be the scalar curvature of a Riemannian metric $g$ on $(M,\partial M)$ and let $A_{g}$ be second fundamental form of $\partial M$. $\partial M$ is totally geodesic if $A_{g}\equiv 0$, totally umbilic if $A_{g}=\lambda g$ for some $\lambda\in C^{\infty}(\partial M),$ minimal if ${\rm Tr}(A_{g})\equiv 0$, and mean convex if ${\rm Tr}(A_{g})$ does not point inward anywhere on the boundary. Let $\eta_{1}$ be the first eigenvalue of the conformal Laplacian $\Box_{g}$ with boundary conditions given in (2). For $M$ a spin $4k$-manifold, let $\xi_{\alpha}(\partial M,g|_{\partial M})$ be the $\mathbb{R}$-valued $\xi$-invariant of $\partial M$ associated to the Dirac operator on $\partial M$, with $\alpha$ a unitary representation of $\pi_{1}(\partial M).$
Theorem 4.1.
Let $(M,\partial M,g)$ be a spin manifold with $\eta_{1}>0$ and totally umbilic boundary. Assume that $\alpha$ is a unitary representation of $\pi_{1}(\partial M)$ that extends to a unitary representation of $\pi_{1}(M).$ Then $\xi_{\alpha}(\partial M,g|_{\partial M})=0.$ In particular, if $M$ admits a metric $g$ with $R_{g}>0$ and $A_{g}\equiv 0$, then $\xi_{\alpha}(\partial M,g|_{\partial M})=0.$
The $\xi$-invariant is “quasi-topological,” in that its $\mathbb{R}/\mathbb{Z}$ reduction is independent of the metric on $\partial M.$ Using this invariant, we give examples of $4$-manifolds not admitting psc metrics with totally geodesic or even mean convex boundary. Let $L(p,q)$ be a lens space in the usual notation.
Theorem 5.1.
Let $p$ be odd. Then there exists a $4$-manifold $M(p,q)$ with boundary $p^{2}L(p,q)$ such that $M(p,q)$ has no metric with $\eta_{1}>0$ and mean convex boundary. In particular, $M(p,q)$ has no psc metric with minimal boundary.
The proof first rules out $\eta_{1}>0$ and totally geodesic boundary, and then appeals to a difficult theorem in [7] to rule out mean convex boundary. We believe these are the first examples in this dimension. The construction of $M(p,q)$ is explicit; bordism theory shows the existence of such a manifold with boundary $pL(p,q)$.
Schematically, Thm. 4.1 is of the form
(1) $R_{g}>0,A_{g}=0\Rightarrow\eta_{1}>0,\zeta_{1}>0\Rightarrow\xi_{\alpha}(\partial M,g|_{\partial M})=0.$
Here $\zeta_{1}$ is the lowest eigenvalue of the conformal Laplacian on $\partial M.$ The first implication follows from the variational definitions of $\eta_{1}$ and $\zeta_{1}$, while the second implication uses [32, Thm. 1.2] and the APS index theorem. Therefore, while the most appealing hypotheses to geometers involve the scalar curvature and the second fundamental form, a conformal analytic assumption like $\eta_{1}>0$ is a better/weaker assumption on the Riemannian metric.
There is a series of implications similar to (1) for closed Riemannian spin manifolds $(M^{4k},g)$. The Lichnerowicz formula and the AS index theorem give the standard obstruction to psc metrics on $M$:
$R_{g}>0\Rightarrow{\rm ker}(\not{\partial}^{\pm}_{g})=0\Rightarrow\hat{A}(M)=0,$
where ∂̸$\not{\partial}_{g}^{\pm}∂$ is the Dirac operator on the sum of plus and minus spinors. This has no immediate connection to conformal geometry. We can insert $\eta_{1}$ somewhat artificially: for $\tilde{g}$ conformally equivalent to $g$ with $R_{\tilde{g}}$ constant,
$R_{g}>0\Rightarrow\eta_{1}>0\ {\rm and}\ \Delta+\frac{R_{\tilde{g}}}{4}>0\Rightarrow{\rm ker}(\not{\partial}_{\tilde{g}}^{\pm})={\rm ker}(\not{\partial}_{g}^{\pm})=0\Rightarrow\hat{A}(M)=0.$
The first implication follows from the solution of the Yamabe conjecture: $\eta_{1}>0$ implies $\tilde{g}$ exists with $R_{\tilde{g}}$ positive. For the second implication, see [26, Thm. 3.5]. In contrast, the use of conformal analytic techniques seems essential in the manifold with boundary case.
For related work, [8] uses $\xi$-invariants to produce examples of psc metrics $g_{1},g_{2}$ on boundary components of a spin manifold $M$ such that $M$ admits no psc metric which is a product metric $dt^{2}\oplus g_{i}$ near the boundary. [2] gives conditions under which a psc metric has a deformation to a psc metric which is a product near the boundary, and [1] discusses when a psc metric on a closed manifold $X$ can be filled in to a psc metric on a manifold $W$ with $\partial W=X$. [11] uses $\xi$-invariants as an obstruction to the filling problem in conformally flat (e.g., hyperbolic) geometry. The metrics in [1, 2, 8] have vanishing second fundamental form and hence totally umbilic and minimal boundary, so the work in §3 can be thought of as the geometric analysis needed to move beyond the product metric case. Using a combination of Ricci flow and other dimension-specific techniques, [9] completely characterizes the topology of all $3$-manifolds with boundary admitting metrics with $R_{g}>0$ and non-negative mean curvature. In a different approach, [7, Thm. 19] gives a differential topological obstruction in terms of $K$-area to a manifold with $R_{g}>0$ and non-negative mean curvature, and produces an example of a $3$-manifold admitting no such metric. In the other direction, the existence of constant psc metrics on manifolds with boundary goes back to [13, 14] (with [14] corrected in [29]), with recent work by the third author [30, 31, 33]. [24, Thms. 5.3, 5.5] gives topological conditions under which a totally non-spin $M$ has a psc metric with vanishing or positive mean curvature. These results have the advantage of having boundary assumptions only on the mean curvature, rather than the second fundamental form.
As an outline of the paper, in §2 we give the analytic setup. As a first conformal analytic/geometric result, we prove that $\eta_{1}>0,\zeta_{1}>0$ implies the existence of a metric $\tilde{g}$ conformal to $g$ with $R_{\tilde{g}},R_{\tilde{g}|_{\partial M}}$ positive (Thm. 2.1). In §3, we relate the positivity of $\Box_{g}$ on $M$ to geometric and analytic conditions on $\partial M.$ In particular, we replace the condition $\zeta_{1}>0$ in Thm. 2.1 with the assumption that $\partial M$ is totally umbilic (Thm. 3.1). We find a geometric obstruction, namely $R_{\bar{g}}\leq 0$, to filling in an $n$-manifold $(X,\bar{g})$ to an $(n+1)$-manifold $(M,g)$ with $\Box_{g}>0$ and $X$ totally umbilic (Cor. 3.2). We similarly find an analytic obstruction to finding a spin manifold $(M,g)$ with $\Box_{g}>0$ and totally umbilic boundary, namely the presence of nontrivial harmonic spinors on $\partial M$ (Cor. 3.4). The transition from the analytic assumption $\zeta_{1}>0$ in §2 to the geometric assumption (totally umbilic boundary) in §3 makes the geometric analysis more difficult: compare the simple boundary argument in Thm. 2.1 with the more difficult proof of Thm. 3.1.
In §4, we prove Thm. 4.1. In §5, we prove Thm. 5.1.
We would like to thank the referee both for pointing out a serious error in an earlier version of the paper and for encouraging us to replace the totally geodesic boundary condition with the weaker totally umbilic condition in §§3,4.
2. Positive Scalar Curvature and Positivity of the conformal Laplacians
Throughout this paper, let $(M^{n},g)$ be a connected, compact, orientable smooth Riemannian manifold with smooth boundary $\partial M$, $n=\dim M\geqslant 3$. In this section, we give analytic conditions, namely the positivity of the conformal Laplacians on $M$ and $\partial M$, under which there exists a metric $\tilde{g}\in[g]$, the conformal class of $g$, such that the scalar curvatures $R_{\tilde{g}}$ on $M$ and $R_{\imath^{*}\tilde{g}}$ on $\partial M$ are positive, for $\imath:\partial M\to M$ the inclusion (Thm. 2.1). In §3, we will replace the analytic boundary condition with geometric boundary conditions (Prop. 3.1, Thm. 3.1), which is then used in §4 (Thm. 4.1).
2.1. Notation and Basic Setup
The unit outward normal vector field along $\partial M$ with metric $g$ is denoted by $\nu=\nu_{g}.$ $R_{g}$ is the scalar curvature of $g$, $\text{Ric}_{g}(X,Y)$ is the Ricci curvature tensor, and $A_{g}$ is the second fundamental form on $\partial M$: $A_{g}(X,Y)=\left(\nabla_{X}Y\right)^{\perp}$ for $X,Y\in T(\partial M)$, where $\nabla$ is the Levi-Civita connection of $g$ and $\perp$ is projection to the normal direction. The mean curvature of $\partial M$ is $h_{g}={\rm Tr}(A_{g})=g\left(\sum_{i}\nabla_{e_{i}}e_{i},\nu\right)$, where $\{e_{i}\}$ is an orthonormal frame of $T_{x}\partial M$ at $x\in\partial M.$
The positive definite Laplace-Beltrami operator is
$-\Delta_{g}u=-\frac{1}{\sqrt{\det(g)}}\partial_{i}\left(g^{ij}\sqrt{\det(g)}\partial_{j}u\right).$
For
$a=\frac{4(n-1)}{n-2},p=\frac{2n}{n-2},$
the conformal Laplacian is
$\Box_{g}u:=-a\Delta_{g}u+R_{g}u;$
other sources define $\Box_{g}$ to be $a^{-1}$ times our $\Box_{g}.$ We use a Robin boundary operator on $\partial M$ as in [13]:
(2) $B_{g}u:=\frac{\partial u}{\partial\nu}+\frac{2}{p-2}h_{g}u.$
$\eta_{1}=\eta_{1,g}$ is the first eigenvalue of conformal Laplacian with respect to the Robin condition $B_{g}u=0$:
(3) $\Box_{g}\varphi=\eta_{1}\varphi\;{\rm in}\;M,B_{g}\varphi=0\;{\rm on}\;\partial M.$
$\zeta_{1}$ is the first eigenvalue of the conformal Laplacian on the closed manifold $\partial M:$ $\Box_{i^{*}g}\psi=\zeta_{1}\psi$ on $\partial M.$
Let $\lambda(\partial M,i^{*}g)$ be the Yamabe invariant on $\partial M$. We have
$\Box_{g}>0\Leftrightarrow\eta_{1}>0,\ \ \Box_{i^{*}g}>0\Leftrightarrow\zeta_{1}>0\Leftrightarrow\lambda(\partial M,i^{*}g)>0.$
Here ${\rm sgn}(\zeta_{1})={\rm sgn}(\lambda(\partial M,i^{*}g))$ by [23, Lemma 2.5], and both signs are conformal invariants.
Let $X$ be a smooth vector field and $f$ a smooth function on $M$. In a local orthonormal frame $\{e_{1},\dotso,e_{n}\}$, the vector field $\nabla_{X}f=X(f)=df(X)$ has the local expression $X^{i}\nabla_{i}f$ for $X=X^{i}e_{i}$. In particular, $\nabla_{n}f=\frac{\partial f}{\partial\nu}$ on $\partial M$.
2.2. The sign of $R_{g}$ and $R_{\imath^{*}g}$
The prescribed scalar curvature problem both in the interior $M$ and on the boundary $\partial M$ is treated in [33]. We recall that for given functions $R\in\mathcal{C}^{\infty}(M)$ and $f\in\mathcal{C}^{\infty}(\partial M)$, the existence of a metric $\tilde{g}\in[g]$ with $R_{\tilde{g}}=R$ and with Dirichlet boundary condition $f$ reduces to the existence of a positive, smooth solution of the Yamabe equation
(4) $\Box_{g}u=Ru^{p-1}\;{\rm in}\;M,u=f\;{\rm on}\;\partial M.$
For an appropriate choice of $f$, we proved that the metric $\tilde{g}=u^{p-2}g=u^{4/(n-2)}g$ has $R_{\tilde{g}}=R$ on $M$ and prescribed scalar curvature $R_{\imath^{*}\tilde{g}}$ on $\partial M$.
Assuming the positivity of both $\eta_{1}$ and $\zeta_{1}$, we can improve the results in [33] to control both the signs of $R_{\tilde{g}}$ and $R_{\imath^{*}\tilde{g}}$ for some $\tilde{g}\in[g]$.
Theorem 2.1.
Assume that $\eta_{1}>0$ for $(M^{n},\partial M,g)$, $n\geqslant 3$. If either (i) $\zeta_{1}>0$ when $n\geqslant 4$, or (ii) $\chi(\partial M_{i})>0$ for each boundary component $\partial M_{i}$ of $\partial M$ when $n=3$, then there exists $\tilde{g}\in[g]$ such that $R_{\tilde{g}}>0$ on $M$ and $R_{\imath^{*}\tilde{g}}>0$ on $\partial M$.
Proof.
Let $\tilde{\eta}_{1}$ be the first eigenvalue of the conformal Laplacian $\Box_{g}$ with vanishing Dirichlet boundary condition, i.e., there exists $\varphi\in C^{\infty}(M)$ with $\Box_{g}\varphi=\tilde{\eta}_{1}\varphi\;{\rm in}\;M,\varphi=0\;{\rm on}\;\partial M.$ By [33, Prop. 4.3], $\tilde{\eta}_{1}\geqslant\eta_{1}>0$. Given $f\in\mathcal{C}^{\infty}(\partial M)$, $f>0$, there exists a positive, smooth function $u$ and a positive constant $\lambda$ such that
(5) $-a\Delta_{g}u+R_{g}u=\lambda u^{p-1}\;{\rm in}\;M,u=f\;{\rm on}\;\partial M,$
by [33, Thm. 4.5].
By the solution of the Yamabe problem on closed manifolds, $\zeta_{1}>0$ implies there exists $\phi\in C^{\infty}(\partial M),\phi>0$, such that the metric $g_{1}=\phi^{\frac{4}{n-3}}\imath^{*}g$ has $R_{g_{1}}>0$ on $\partial M$, provided $n\geqslant 4$. When $n=3$, there exists a conformal change $g_{1}=e^{2\varphi}\imath^{*}g$ with Gaussian curvature $R_{g_{1}}>0$ on $\partial M$ [23].
Set
(6) $f=\begin{cases}\left(\phi^{\frac{4}{n-3}}\right)^{\frac{n-2}{4}},&n\geqslant 4,\\ \left(e^{2\varphi}\right)^{\frac{n-2}{4}},&n=3.\end{cases}$
Using this $f$ in (5), we conclude that $\tilde{g}=u^{p-2}g$ has $R_{\tilde{g}}>0$ on $M$ and $R_{\imath^{*}\tilde{g}}>0$ on $\partial M$. ∎
3. Positivity of the conformal Laplacian: Geometric and Analytic boundary obstructions
A Riemannian manifold with psc metric $g$ and minimal boundary has positive conformal Laplacian: $\eta_{1}>0$. In this section, we conversely assume $\eta_{1}>0$ and discuss geometric implications. For example, in Cor. 3.1 we give a proof that $\eta_{1}>0$ implies the existence of a psc $\tilde{g}\in[g]$ with minimal boundary, a type of conformal converse. If we extend the hypotheses to $\eta_{1}>0$ and $\partial M$ is totally geodesic, then we find $\tilde{g}\in[g]$ with $R_{\imath^{*}\tilde{g}}>0$ and $\partial M$ still totally geodesic (Prop. 3.1). With more work, we can weaken the boundary condition to totally umbilic and find $\tilde{g}\in[g]$ with $R_{\tilde{g}}$ and $R_{\imath^{*}\tilde{g}}$ positive (Thm. 3.1).
These results can be stated as boundary geometric obstructions to $\eta_{1}>0$. For example, Cor. 3.2 states that a closed manifold $(X,\bar{g})$ with $R_{\bar{g}}\leq 0$ cannot be the boundary of a manifold $(M,g)$ with $\eta_{1}>0$ and $\imath^{*}g$ conformal to $\bar{g}.$ We also obtain a boundary analytic obstruction in Cor. 3.4: if a spin manifold has $\eta_{1}>0$ and totally geodesic boundary, then the boundary has no harmonic spinors.
The proofs of these results are much simpler on the class of metrics which are products near the boundary. Thus this section highlights the work necessary to pass first from the product metric case to the totally geodesic case and then to the totally umbilic case.
The motivation for assuming totally geodesic boundary (rather than minimal boundary) is the following. The relation between $R_{g}$ and $R_{\imath^{*}g}$ is given by taking the complete trace of the Gauss-Codazzi equation:
(7) $R_{g}=2\text{Ric}_{g}(\nu_{g},\nu_{g})+R_{\imath^{*}g}+\lvert A_{g}\rvert^{2}-h_{g}^{2}\;{\rm on}\;\partial M.$
The vanishing of the second fundamental form reduces this equation to the more manageable
(8) $R_{g}=2\text{Ric}_{g}(\nu,\nu)+R_{\imath^{*}g}.$
The analysis of this equation under conformal change is key to the results in this section. With extra work, we can weaken the boundary condition to totally umbilic; see §3.2.
3.1. Preliminary results on positivity with totally geodesic or minimal boundary
We collect some known or folklore results about positivity of scalar curvature if $\eta_{1}>0$ and the boundary is minimal or totally geodesic. Proofs are included for completeness.
Remark 3.1.
We first note that Neumann boundary conditions are favorable in conformal geometry. For $\tilde{g}=e^{2\phi}g$, the second fundamental form and mean curvature transform by
(9) $A_{\tilde{g}}(X,Y)$ $=e^{\phi}A_{g}(X,Y)+\frac{\partial\phi}{\partial\nu}e^{\phi}g(X,Y),$
$h_{\tilde{g}}$ $=e^{-\phi}h_{g}+\frac{\partial\phi}{\partial\nu}e^{-\phi},$
where both sides are multiples of the outward unit normal vector for $g$. Thus under this conformal change with Neumann boundary condition $\frac{\partial\phi}{\partial\nu}=0$, the vanishing of the second fundamental form or mean curvature for $g$ implies the same for $\tilde{g}$. Note that $\frac{\partial\phi}{\partial\nu}=0$ is equivalent to $\frac{\partial u}{\partial\nu}=0$ for
(10) $\tilde{g}=u^{p-2}g=e^{2\phi}g.$
Assuming $\eta_{1}>0$ and good boundary geometry, we can alter a metric within its conformal class to have positive scalar curvature.
Proposition 3.1.
If $(M^{n},\partial M,g)$ has $\eta_{1}>0$ and $\partial M$ totally geodesic, resp. minimal with $n\geqslant 3$, then there exists a metric $\tilde{g}$ conformal to $g$ such that $R_{\tilde{g}}>0$ and $\partial M$ remains totally geodesic, resp. minimal, for $\tilde{g}$. If $\partial M$ is totally geodesic, then $R_{i^{*}\tilde{g}}>0.$
Proof.
We treat the case where $\partial M$ is totally geodesic; the minimal boundary case is similar. By [13, §1], the eigenvalue problem for the conformal Laplacian
(11) $-a\Delta_{g}\varphi+R_{g}\varphi=\eta_{1}\varphi\;{\rm in}\;M,\frac{\partial\varphi}{\partial\nu}=0\;{\rm on}\;\partial M,$
admits a smooth, positive solution $\varphi$. By the maximum principle, $\varphi>0$ on $M$, including on the boundary. The mean curvature term $h_{g}$ in the Robin boundary condition disappears, since $A_{g}\equiv 0$. Rewriting (11) as
(12) $-a\Delta_{g}\varphi+R_{g}\varphi=(\eta_{1}\varphi^{2-p})\varphi^{p-1}\;{\rm in}\;M,\frac{\partial\varphi}{\partial\nu}=0\;{\rm on}\;\partial M,$
it follows from (12) that the metric $\tilde{g}=\varphi^{p-2}g$ has scalar curvature $R_{\tilde{g}}=\eta_{1}\varphi^{2-p}>0$. For $\tilde{g}=\varphi^{p-2}g:=e^{\psi}g$, the boundary condition in (12) gives $\frac{\partial\psi}{\partial\nu}\equiv 0\;{\rm on}\;\partial M.$ Since $A_{g}\equiv 0$, (9) yields $A_{\tilde{g}}(X,Y)=0.$
For the last statement, it is standard that a collar of $\partial M$ has coordinates $(x,t)\in\partial M\times[0,1)$ such that $\tilde{g}(x,t)=h(x,t)\oplus dt^{2}$, for some family of metrics $h(x,t)$ on $\partial M$. If $\partial M$ is totally geodesic, a straightforward calculation in these coordinates gives $R_{\tilde{g}}|_{\partial M}=R_{i^{*}\tilde{g}}.$
Similarly, we can reprove the known result that $\eta_{1}>0$ implies the positivity of scalar curvature and vanishing of mean curvature for some $\tilde{g}\in[g].$
Corollary 3.1.
If $(M^{n},\partial M,g)$ has $\eta_{1}>0$, then there exists a metric $\tilde{g}\in[g]$ such that $R_{\tilde{g}}>0$ and $h_{\tilde{g}}=0$.
Proof.
Again by [13, §1], the eigenvalue problem for the conformal Laplacian
$-a\Delta_{g}\varphi+R_{g}\varphi=\eta_{1}\varphi=(\eta_{1}\varphi^{2-p})\varphi^{p-1}\;{\rm in}\;M,\frac{\partial\varphi}{\partial\nu}+\frac{2}{p-2}h_{g}=0\;{\rm on}\;\partial M,$
admits a smooth, positive solution $\varphi$. As in (12), the metric $\tilde{g}=\varphi^{p-2}g$ has the desired properties. ∎
3.2. Positivity with totally umbilic boundary
In this subsection, we weaken the totally geodesic boundary assumption in Prop. 3.1 to totally umbilic (Thm. 3.1). The main point is that we can still guarantee that $R_{i^{*}\tilde{g}}>0$ on $\partial M.$
Definition 3.1.
Let $(M^{n},\partial M,g)$ be a compact Riemannian manifold with boundary. $\partial M$ is totally umbilic if the second fundamental form $A_{g}$ is proportional to the first fundamental form $g$, i.e.
(13) $A_{g}(X,Y)=\lambda_{p}g(X,Y),\forall X,Y\in T_{p}\partial M,p\in\partial M.$
Considering $A_{g}$ as a linear transformation on $T_{p}\partial M$, $\partial M$ is totally umbilic iff all eigenvalues of $A_{g}$ are equal to $\lambda_{p}$. The mean curvature is then
(14) $h_{g}(p)=(n-1)\lambda_{p}.$
From this, it is immediate that a boundary is totally geodesic iff it is minimal and totally umbilic. The Clifford torus is a minimal submanifold of $S^{3}$ which is not totally umbilic, and shows that the positivity of scalar curvature on a manifold with minimal boundary does not imply positivity of the scalar curvature on the boundary. [10, Ex. 4.4] gives an example of a totally umbilic submanifold which is not minimial.
Being totally umbilic is a conformally invariant condition.
Proposition 3.2.
[11, Prop. 2.1] The multiplicities of the eigenvalues of $A_{g}$ are conformal invariants. In particular, $g$ is totally umbilic iff $\tilde{g}\in[g]$ is totally umbilic.
We first relate $R_{g}$ to $R_{i*g}$ in the totally umbilic case. Choose an othonormal local frame $e_{1},\dotso,e_{n}=\nu_{g}$ at a fixed point $p\in\partial M.$ Applying (13 and (14) at $p$, we have
$h_{g}^{2}(p)-\lvert A_{g}\rvert(p)^{2}=\left((n-1)\lambda_{p}\right)^{2}-\sum_{i,j=1}^{n-1}\left(\lambda_{p}g(e_{i},e_{j})\right)^{2}=(n-1)(n-2)\lambda_{p}^{2}\geqslant 0.$
By Gauss-Codazzi (7),
(15) $R_{\imath^{*}g}(p)=R_{g}(p)-2\text{Ric}_{g}(\nu_{g},\nu_{g})(p)+h_{g}^{2}(p)-\lvert A_{g}\rvert(p)^{2}\geqslant R_{g}(p)-2\text{Ric}_{g}(\nu_{g},\nu_{g})(p).$
In particular, $R_{g}-2\text{Ric}_{g}(\nu_{g},\nu_{g})>0$ on $\partial M$ implies $R_{\imath^{*}g}>0$ on $\partial M$.
We therefore focus on the sign of $R_{\tilde{g}}-2\text{Ric}_{\tilde{g}}(\nu_{\tilde{g}},\nu_{\tilde{g}})$ on $\partial M$ for some metric $\tilde{g}\in[g]$. Set $\text{Ric}_{g,nn}=\text{Ric}_{g}(\nu_{g},\nu_{g})$, and similarly for $\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}.$ In terms of $\phi$, the well-known transformation laws for the Ricci curvature tensor and the scalar curvature are
(16) $\begin{split}\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}&=e^{-2\phi}\left(\text{Ric}_{g,nn}-(n-2)\left(\nabla_{n}\nabla_{n}\phi-\nabla_{n}\phi\nabla_{n}\phi\right)-\left(\Delta_{g}\phi+(n-2)\lvert d\phi\rvert_{g}^{2}\right)g_{nn}\right);\\ R_{\tilde{g}}&=e^{-2\phi}\left(R_{g}-2(n-1)\Delta_{g}\phi-(n-2)(n-1)\lvert d\phi\rvert_{g}^{2}\right).\end{split}$
Using (16), we have
$-2\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}+R_{\tilde{g}}$ $=e^{-2\phi}\left(-2\text{Ric}_{g,nn}+R_{g}-2(n-1)\Delta_{g}\phi+2\Delta_{g}\phi g_{nn}+2(n-2)\nabla_{n}\nabla_{n}\phi\right)$
$\qquad+e^{-2\phi}\left(-2(n-2)\nabla_{n}\phi\nabla_{n}\phi+2(n-2)\lvert d\phi\rvert_{g}^{2}g_{nn}-(n-2)(n-1)\lvert d\phi\rvert_{g}^{2}\right).$
Rewriting this with respect to both conformal factors $u,\phi$, we have
${\left(-2\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}+R_{\tilde{g}}\right)|_{\partial M}}$
$=e^{-2\phi}\left(-2\text{Ric}_{g,nn}+2(n-2)\left(\nabla_{n}\nabla_{n}\phi-\nabla_{n}\phi\nabla_{n}\phi\right)+\left(2\Delta_{g}\phi+2(n-2)\lvert d\phi\rvert_{g}^{2}\right)g_{nn}\right)$
$\qquad+e^{-2\phi}\left(R_{g}-2(n-1)\Delta_{g}\phi-(n-1)(n-2)\lvert d\phi\rvert_{g}^{2}\right)$
$=e^{-2\phi}\left(-2\text{Ric}_{g,nn}+4u^{-1}\nabla_{n}\nabla_{n}u-3(n-2)\nabla_{n}\phi\nabla_{n}\phi\right)$
(17) $\qquad+e^{-2\phi}\left(u^{-1}\cdot\frac{4}{n-2}\Delta_{g}u+(n-2)\lvert d\phi\rvert_{g}^{2}\right)\cdot g_{nn}$
$\qquad\qquad+e^{-2\phi}\left(R_{g}-u^{-1}\frac{4(n-1)}{n-2}\cdot\Delta_{g}u\right)$
$=e^{-2\phi}u^{-1}\cdot\frac{1}{n-2}\left(-(4n-4-4g_{nn})\Delta_{g}u+(4n-8)\nabla_{n}\nabla_{n}u+(n-2)R_{g}u\right)$
$\qquad+e^{-2\phi}u^{-1}\cdot\frac{1}{n-2}\left(-2(n-2)u\text{Ric}_{g,nn}+(n-2)^{2}u\lvert d\phi\rvert_{g}^{2}g_{nn}-3(n-2)^{2}u\nabla_{n}\phi\nabla_{n}\phi\right).$
The next step is to prove the existence of a smooth solution of the following PDE, which contains the linear terms in the second to the last line of (3.2), i.e.,
$-(4n-4-4g_{nn})\Delta_{g}u+(4n-8)\nabla_{n}\nabla_{n}u+(n-2)R_{g}u=F\;{\rm in}\;M,\frac{\partial u}{\partial\nu}=0\;{\rm on}\;\partial M,$
for some appropriate choice of smooth function $F$.
Under our standing assumption $\eta_{1}>0$, we may assume $R_{g}>0$ by Theorem 3.1 (where $\tilde{g}$ in that result is now relabeled $g$). We may also assume $g_{nn}<1$ on $\partial M$ by appropriate choices of local coordinates. In other words, for coordinates $(x^{1},\ldots,x^{n})$ near $p\in\partial M$ with $\nu_{p}=\partial/\partial x^{n}$, define new coordinates $(y^{1}=x^{1},\ldots,y^{n-1}=x^{n-1},y^{n}=\alpha^{-1}x^{n})$ for a constant $\alpha<1.$ Since $\nu$ is globally defined, in the $y$ coordinates $g_{nn}=\alpha^{2}$ on all of $\partial M.$ (16) and (3.2) still hold in the $y$ coordinates.
Let $C$ be a closed collar of $\partial M$, so $C$ is diffeomorphic to $\partial M\times[0,1]$, with $\partial M=\partial M\times\{0\}.$
Lemma 3.1.
For $F\in\mathcal{C}^{\infty}(M)$, the partial differential equation
(18) $-(4n-4-4g_{nn})\Delta_{g}u+(4n-8)\nabla_{n}\nabla_{n}u+(n-2)R_{g}u=F\;{\rm in}\;C,\frac{\partial u}{\partial\nu}=0\;{\rm on}\;\partial C$
admits a solution $u\in\mathcal{C}^{\infty}(C)$.
Proof.
Since $g_{nn}<1$, we claim that the differential operator is uniformly elliptic on $M$. Indeed, the second order symbol at a cotangent vector $\xi=\xi_{i}dy^{i}$ equals
$(4n-4-4g_{nn})$ $|\xi|^{2}-(4n-8)\xi_{n}^{2}=(4n-4-4g_{nn})g^{ij}\xi_{i}\xi_{j}-(4n-8)\xi_{n}^{2}$
$>(4n-8)|\xi^{\perp}|^{2}+(4n-8)\xi_{n}^{2}-(4n-8)\xi_{n}^{2}\geq 0,$
where $\xi^{\perp}:=\xi-\xi_{n}dy^{n},$ and we use $g^{in}=0$ for $i\neq n$. In addition, the zeroth order term $(n-2)R_{g}$ is positive, so the differential operator in (18) is injective. Applying the standard Lax-Milgram argument, we conclude that there exists $u\in H^{1}(M,g)$ solving (18). The smoothness of $u$ follows from standard elliptic regularity arguments. ∎
We now investigate the relationship between $R_{\tilde{g}}$ on $\partial M$ and $R_{i^{*}\tilde{g}}$ under the conformal change $\tilde{g}=u^{p-2}g$, where $u$ is obtained by first solving (18) on $C$ for an appropriate $F$, and then extending the solution to a positive function $u$ on $M$ by partition of unity. Of course, $u$ may not solve (18) outside the collar region.
Choose $F\in C^{\infty}(\partial M)$ such that $F-2(n-2){\rm Ric}_{g,nn}>0\;{\rm on}\;\partial M.$ Fix an integer $q\gg 1$ such that the Sobolev embedding $W^{2,q}(C,g)\hookrightarrow\mathcal{C}^{1,\alpha}(C)$ is continuous. Extend $F$ smoothly to all of $C$ such that $\lVert F\rVert_{\mathcal{L}^{q}(C,g)}\ll 1$, i.e. the $\mathcal{L}^{q}$-norm is small enough, where this size will be determined later. So far, we have
(19) $\lVert F\rVert_{\mathcal{L}^{q}(C,g)}\ll 1,F-2(n-2)Ric_{g,nn}>0\;{\rm on}\;\partial M.$
By Lemma 3.1, there exists $u_{0}\in C^{\infty}(C)$ that solves (18).
We now show that $u_{0}$ is $\mathcal{C}^{0}$-small for our choice of $F$. The standard $\mathcal{L}^{q}$-type elliptic regularity estimate on $(C,\partial C,g)$ is
(20) $\lVert u_{0}\rVert_{W^{2,q}(C,g)}\leqslant C_{0}\left(\lVert Lu_{0}\rVert_{\mathcal{L}^{q}(C,g)}+\lVert u_{0}\rVert_{\mathcal{L}^{q}(C,g)}\right).$
for a second order uniformly elliptic operator $L$, provided that $u_{0}\in H^{1}(C,g)\cap L^{q}(C,g)$. The constant $C_{0}$ is independent of $u_{0}$. When $L:W^{2,q}(C,g)\rightarrow\mathcal{L}^{q}(C,g)$ is injective, as in Lemma 3.1, we can improve (20) to
(21) $\lVert u_{0}\rVert_{W^{2,q}(C,g)}\leqslant C_{0}\lVert Lu_{0}\rVert_{\mathcal{L}^{q}(C,g)}.$
For $q$ as above, the Sobolev embedding yields
(22) $\lVert u_{0}\rVert_{\mathcal{C}^{0}}\leqslant C^{\prime}\lVert u_{0}\rVert_{W^{2,q}(C,g)}.$
Applying (21) and (22) to the solution $u_{0}$ of (18), we have
(23) $\lVert u_{0}\rVert_{\mathcal{C}^{0}}\leqslant C^{\prime}C_{0}\lVert F\rVert_{\mathcal{L}^{q}(C,g)}.$
We now choose $\lVert F\rVert_{\mathcal{L}^{q}(C,g)}$ in (19) small enough so that $\lVert u_{0}\rVert_{\mathcal{C}^{0}}\ll 1$.
The conformal factor $u$ needs to be positive. For $D:=\lVert u_{0}\rVert_{\mathcal{C}^{0}}$, set
(24) $u_{1}:=u_{0}+(1+\epsilon)D$
for some $\epsilon\ll 1$. It is clear that $u_{1}>0$. By modifying $F$ if necessary, we can take $u_{0}$ small enough so that $u_{1}<1$. By (18),
(25) $-(4n-4-4g_{nn})\Delta_{g}u_{1}$ $+(4n-8)\nabla_{n}\nabla_{n}u_{1}+(n-2)R_{g}u_{1}=F+(n-2)(1+\epsilon)R_{g}D\;{\rm in}\;C,$
$\frac{\partial u_{1}}{\partial\nu}$ $=0\;{\rm on}\;\partial M.$
Taking the conformal transformation $\tilde{g}=u_{1}^{p-2}g:=e^{\phi_{1}}g,$ it follows from (3.2) and (25) that
${\left(-2\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}+R_{\tilde{g}}\right)|_{\partial M}}$
$=e^{-2\phi_{1}}u_{1}^{-1}\cdot\frac{1}{n-2}\left(-(4n-4-4g_{nn})\Delta_{g}u_{1}+(4n-8)\nabla_{n}\nabla_{n}u_{1}+(n-2)R_{g}u_{1}\right)|_{\partial M}$
$\qquad+e^{-2\phi_{1}}u_{1}^{-1}\cdot\frac{1}{n-2}\left(-2(n-2)u_{1}\text{Ric}_{g,nn}+(n-2)^{2}u_{1}\lvert d\phi_{1}\rvert_{g}^{2}g_{nn}\right)$
$\qquad-e^{-2\phi_{1}}u_{1}^{-1}\cdot\frac{1}{n-2}\left(3(n-2)^{2}u_{1}\nabla_{n}\phi_{1}\nabla_{n}\phi_{1}\right)|_{\partial M}$
$\geqslant e^{-2\phi_{1}}u_{1}^{-1}\cdot\frac{1}{n-2}\left(F+(n-2)(1+\epsilon)DR_{g}-2(n-2)u_{1}\text{Ric}_{g,nn}\right)$
$>0,$
where the last inequality uses (19) and $10<u_{1}<$ 1.
We now extend $u_{1}$ from $C$ to $M$ so that $u_{1}$ is positive on all of $M$. $u_{1}$ satisfies (18) near $\partial M$. By (15), we conclude the following:
Theorem 3.1.
If $(M^{n},\partial M,g)$ has $\eta_{1}>0$ and $\partial M$ totally umbilic, then there exists a metric $\tilde{g}$ conformal to $g$ such that $R_{\tilde{g}}>0$ and $R_{\imath^{*}\tilde{g}}>0$.
This result gives a conformal geometric obstruction on $\partial M$ to the positivity of $\Box_{g}$ on $M$, and can be thought of as an obstruction to filling.
Corollary 3.2.
Let $(X^{n},\bar{g})$ be a closed Riemannian manifold, with $n\geqslant 2$. Let $\zeta_{1}=\zeta_{1,\bar{g}}$ be the first eigenvalue of the conformal Laplacian $\Box_{\bar{g}}$ on $X$.
  1. (i)
    If $n\geqslant 3$ and $R_{\bar{g}}\leqslant 0$ everywhere on $X$, there exists no compact manifold $(M,g)$ with $\eta_{1}>0$ and totally umbilic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$.
  2. (ii)
    If $n\geqslant 3$ and $\zeta_{1}\leqslant 0$ on $X$, there exists no compact manifold $(M,g)$ with $\eta_{1}>0$ and totally umbilic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$.
  3. (iii)
    If $n=2$ and $\chi(X_{i})\leqslant 0$ for any component $X_{i}$ of $X$, there is no compact $3$-manifold $(M,g)$ with $\eta_{1}>0$ and totally umbilic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$; similarly, there is no $3$-manifold $(M,g)$ with $\eta_{1}>0$, such that $\partial M=X$, $i^{*}g$ conformal to $\bar{g}$, and $\bar{g}$ has nonpositive Yamabe invariant on any $X_{i}.$
Proof.
For (i), assume such an $M$ exists. $R_{\bar{g}}\leq 0$ implies $\zeta_{1,\bar{g}}\leq 0.$ By the hypotheses and Thm. 3.1, there exists $\tilde{g}$ conformal to $g$ with $R_{\imath^{*}\tilde{g}}>0.$ Thus $\zeta_{1,\imath^{*}\tilde{g}}>0.$ However, $\imath^{*}\tilde{g}$ is conformal to $\bar{g}$ by hypothesis, so the sign of $\zeta_{1}$ for these metrics must be the same.
(ii) follows from the proof of (i).
For (iii), if $g$ exists, there exists $\tilde{g}$ conformal to $g$ with $R_{\imath^{*}\tilde{g}}>0$ by Thm. 3.1. This implies $\chi(X_{i})>0$ for all $i$. $R_{\imath^{*}\tilde{g}}>0$ also implies positive Yamabe invariant on each $X_{i}.$
Example 3.1.
To restate Cor. 3.2(i) geometrically, if $R_{\bar{g}}\leqslant 0$ on $X$, then there exists no compact manifold $(M,g)$ with $R_{g}>0$ and $\partial M$ totally umbilic, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$. This is easy if $g$ is a product metric near $\partial M$: $R_{i^{*}g}=R_{g}>0$ on $\partial M$, so $\zeta_{1,i^{*}g}\geqslant 0.$ On the other hand, $R_{\bar{g}}\leqslant 0$ implies $\zeta_{1,\bar{g}}\leqslant 0.$ As in the previous proof, this is impossible. The amount of analysis in §3 indicates the large gap between the product and nonproduct cases.
Remark 3.2.
An interesting contrast to Cor. 3.2(i) is in [19, Lem. 2.4], where they prove that if $(X,\bar{g})$ has positive Yamabe invariant and nonvanishing conformal invariant $I$, then no $M$ exists with a totally geodesic psc metric conformal to $\bar{g}$ on $X$. We obtain the same conclusion (with totally umbilic boundary) under the hypothesis of nonpositive Yamabe invariant.
We now find an analytic obstruction on $\partial M$, assumed to be spin, to the condition $\eta_{1}>0,$ using the following result:
Theorem 3.2.
[33, Thm. 1.1] On $(M^{n},\partial M,g)$, there exists a metric $\tilde{g}$ conformal to $g$ with $R_{\tilde{g}}$ constant on the interior of $M$ and $R_{i^{*}\tilde{g}}$ constant on $\partial M.$
When $\eta_{1}>0$, Thm. 3.2 implies that $R_{\tilde{g}}$ is a positive constant. With the help of Thm. 3.1, we can determine that $R_{i^{*}\tilde{g}}$ is also a positive constant. This is an improvement of Thm. 2.1.
Corollary 3.3.
Let $(M^{n},\partial M,g)$ have totally umbilic boundary. If $\eta_{1}>0$, there exists a metric $\tilde{g}\in[g]$ with $R_{\tilde{g}}$ a positive constant on the interior of $M$ and $R_{i^{*}\tilde{g}}$ a positive constant on $\partial M.$
Proof.
By Thm. 3.1, $\zeta_{1}>0$ when $n>3$ and $\chi(\partial M_{i})>0$ when $n=3$. The rest of the proof follows exactly as in Thm. 2.1 (with $R_{g_{1}}$ a positive constant). ∎
This Corollary gives the analytic obstruction on $\partial M$ mentioned above.
Corollary 3.4.
Let $(M^{n},\partial M,g)$ have $\partial M$ totally umbilic and spin. If $\eta_{1}>0$, then $(\partial M,i^{*}g)$ has no nontrivial harmonic spinors.
Of course, as a boundary $\partial M$ has vanishing $\hat{A}$-genus, so this obstruction is not topological.
Proof.
For $\tilde{g}$ as in Cor. 3.3, $R_{i^{*}\tilde{g}}$ is a positive constant. Now we apply the usual vanishing argument: by the Lichnerowicz formula, the total Dirac operator ∂̸$\not{\partial}_{\partial M}^{\pm}∂$ satisfies
∂̸$\not{\partial}_{\partial M}^{\pm,*}\not{\partial}_{\partial M}^{\pm}=\nabla^{*}\nabla+\frac{1}{4}R_{i^{*}\tilde{g}},∂$
where $\nabla$ is the spinor connection associated to $i^{*}\tilde{g}$. This implies that there are no harmonic spinors for $i^{*}\tilde{g}$. The conclusion for $i^{*}g$ follows, since the dimension of the space of harmonic spinors is a conformal invariant. ∎
Example 3.2.
[…]
4. Positivity of the conformal Laplacian: A quasi-topological obstruction
In this section, we find a “quasi-topological” obstruction to a spin manifold with boundary admitting a metric with $\eta_{1}>0$, namely an $\mathbb{R}$-valued $\xi$-invariant whose reduction mod $\mathbb{Z}$ is topological. As usual, this is also an obstruction to psc metrics with minimal or constant positive mean curvature on the boundary.
Let $(M^{n},g),n=4k$, be a spin manifold with boundary $\partial M$, which is automatically spin. Let $S=S^{+}\oplus S^{-}$ be the spinor bundle on $M$, and let ∂̸=∂̸$\not{\partial}_{g}:\Gamma(S^{+})\to\Gamma(S^{-})∂̸ = ∂$ be the Dirac operator on positive spinors on $M$. Let $(E=E_{\alpha},\nabla)$ be a flat bundle with connection associated to a unitary representation $\alpha$ of $\pi_{1}(M)$. The coupled Dirac operator is ∂̸$\not{\partial}^{E}_{g}=\not{\partial}\otimes{\rm Id}+{\rm Id}\otimes\nabla:\Gamma(S^{+}\otimes E)\to\Gamma(S^{-}\otimes E)∂$.
If $M$ has a metric $g$ with $\eta_{1}>0$, then we can conformally change $g$ to $\tilde{g}$ so that $R_{\tilde{g}}>0$ and $h_{\tilde{g}}=0$ by Corollary 3.1.
Lemma 4.1.
Let $(M,g)$ have $\eta_{1}>0$, and let $\tilde{g}$ be a conformally equivalent metric with $R_{\tilde{g}}>0$ and $h_{\tilde{g}}=0.$ Then $\operatorname{ind}(\not{\partial}_{\tilde{g}})=0$ for the Dirac operator with APS boundary conditions, and $\operatorname{ind}(\not{\partial}^{E}_{\tilde{g}})=0$ for the coupled Dirac operator with APS boundary conditions.
Proof.
The index $\operatorname{ind}(\not{\partial}_{\tilde{g}})$ of the Dirac operator with APS boundary conditions vanishes by [21, Thm. 4]: If $h_{\tilde{g}}\geqslant 0$, then the smallest eigenvalue $\lambda_{1}$ of ∂̸$\not{\partial}^{E}∂$ with APS boundary conditions satisfies
$\lambda_{1}^{2}\geqslant\frac{n+1}{4n}\min_{M}R_{\tilde{g}}.$
Since $R_{\tilde{g}}>0$, we cannot have $\lambda_{1}=0$.
For the coupled Dirac operator, we just follow the proof of this theorem, replacing spinors $\psi\in\Gamma(S)$ by sections $\psi\otimes s:=a^{ij}\psi_{j}\otimes s_{i}\in\Gamma(S\otimes E)$. We can assume $\{\psi_{j}\},\{s_{i}\}$ are local orthonormal parallel sections of $S,E$, resp., at a fixed center point $x$ of a coordinate chart $U$, with $a^{ij}\in C^{\infty}(U),$ by taking synchronous coordinates: take orthonormal frames of $S$ and $E$ at $x$, and parallel translate them out (e.g., geodesic) rays starting at $x$ and filling out $U$. In addition, we assume that $\{\psi_{j}\}$ is constructed in the standard way from a synchronous orthonormal frame $\{e_{k}\}$. Specifically, we need equations (1) - (17) and (the $\geqslant$ part of) Thm. 4 in [21] for $\psi\otimes s.$
At $x$, the key simplifications are (dropping the $\tilde{g}$)
∂̸$\not{\partial}^{E}(\psi\otimes s)∂$ =∂̸$\not{\partial}(a^{ij}\psi_{j})\otimes s_{i}=e_{k}(a^{ij})e_{k}\cdot\psi_{j}\otimes s_{i},= ∂$
$\nabla_{X}^{S\otimes E}(\psi\otimes s)$ $:=(\nabla_{X}^{S}\otimes{\rm Id}+{\rm Id}\otimes\nabla_{X}^{E})(a^{ij}\psi_{j}\otimes s_{i})=\nabla_{X}^{S}(a^{ij}\psi_{j})\otimes s_{i}=X(a^{ij})\cdot\psi_{j}\otimes s_{i}.$
Here the dot is Clifford multiplication, and we use ∂̸$\not{\partial}\psi_{j}=e_{k}\nabla^{S}_{e_{k}}\psi_{j}=0∂$ and $\nabla^{E}s_{i}=0$ at $x$. For example, for $X\in TM$,
$X\langle\psi\otimes s,\phi\otimes s^{\prime}\rangle$ $=X\left(\langle a^{ij}\psi_{j},b^{kl}\psi_{l}\rangle_{S}\langle s_{i},s_{k}\rangle_{E}\right)=X\langle a^{ij}\psi_{j},b^{il}\psi_{l}\rangle_{S}$
$=\sum_{i}\langle\nabla_{X}^{S}(a^{ij}\psi_{j}),b^{il}\psi_{l}\rangle+\sum_{i}\langle a^{ij}\psi_{j},\nabla_{X}^{S}(b^{il}\psi_{l})\rangle$
$=\langle\nabla_{X}^{S\otimes E}(\psi\otimes s),\phi\otimes s^{\prime}\rangle+\langle\psi\otimes s,\nabla_{X}^{S\otimes E}(\phi\otimes s^{\prime})\rangle,$
where the second line uses [21, (1)]: $X\langle\psi,\phi\rangle_{S}=\langle\nabla^{S}_{X}\psi,\phi\rangle+\langle\psi,\nabla_{X}^{S}\phi\rangle.$ Thus we obtain the twisted version of [21, (1)] from the untwisted version.
Similarly, we get the twisted Lichnerowicz formula
(26) (∂̸$\not{\partial}^{E})^{*}\not{\partial}^{E}=(\nabla^{S\otimes E})^{*}\nabla^{S\otimes E}+\frac{R}{4}\otimes{\rm Id}=\left((\nabla^{S})^{*}\nabla^{S}+\frac{R}{4}\right)\otimes{\rm Id}( ∂$
from the Lichnerowicz formula ∂̸∂̸$\not{\partial}^{*}\not{\partial}=(\nabla^{S})^{*}\nabla^{S}+\frac{R}{4}.∂$ As a final example, in our notation [21, (9)] is
(27) ∂̸$\not{\partial}^{S|_{\partial M}}\psi=\frac{n}{2}h\psi-\nu\cdot\not{\partial}^{S}\psi-\nabla^{S}_{\nu}\psi,∂$
where $\nu$ is the unit inward normal vector on $\partial M$. Then by (27),
∂̸$\not{\partial}^{(S\otimes E)|_{\partial M}}(\psi\otimes s)∂$ =∂̸$\not{\partial}^{(S\otimes E)|_{\partial M}}(a^{ij}\psi_{j}\otimes s_{i})=\not{\partial}^{S|_{\partial M}}(a^{ij}\psi_{j})\otimes s_{i}= ∂$
$=\left(\frac{n}{2}ha^{ij}\psi_{j}-\nu\cdot\not{\partial}^{S}(a^{ij}\psi_{j})-\nabla^{S}_{\nu}(a^{ij}\psi_{j})\right)\otimes s_{i}$
$=\frac{n}{2}h(\psi\otimes s)-\nu\cdot\not{\partial}^{S\otimes E}(\psi\otimes s)-\nabla^{S\otimes E}_{\nu}(\psi\otimes s).$
This is the twisted version of [21, (9)].
A straightforward but long check yields the twisted versions of (1)-(17), cumulating in the twisted version of [21, Thm. 4]: we still have $\lambda_{1}^{2}\geqslant[(n+1)/4n]\cdot\min_{M}R_{\tilde{g}}.$ Thus the index of ∂̸$\not{\partial}^{E}_{\tilde{g}}∂$ vanishes. ∎
We next show that the index of the Dirac operator with APS boundary conditions is a conformal invariant, whether or not $\eta_{1}>0.$
Lemma 4.2.
For $\tilde{g}$ conformal to $g$, $\operatorname{ind}(\not{\partial}_{g})=\operatorname{ind}(\not{\partial}_{\tilde{g}}),$ and $\operatorname{ind}(\not{\partial}^{E}_{g})=\operatorname{ind}(\not{\partial}^{E}_{\tilde{g}}).$
Proof.
The APS theorem for manifolds with non-product metric near the boundary [15, §3] is
(28) $\operatorname{ind}(\not{\partial}_{g})=\int_{M}\hat{A}(\Omega_{g})+\int_{\partial M}P(g)-\frac{1}{2}({\rm dim\ ker}(\not{\partial}_{i^{*}g})+\eta_{g}(0)),$
where $\hat{A}(\Omega_{g})$ is the $\hat{A}$-polynomial in the curvature of $g$, and $P(g)$ is the Chern-Simons form related to the $\hat{A}$-polynomial. As a polynomial in the Pontrjagin classes, $\hat{A}(\Omega_{g})$ is a conformal invariant. ${\rm Dim\ ker}(\not{\partial}_{i^{*}g})$ is a conformal invariant, since the Dirac operator is conformally covariant, i.e.. for $\tilde{g}=e^{2\phi}g$, we have ∂̸$\not{\partial}_{\tilde{g}}=e^{-(n+1)\phi/2}\not{\partial}_{g}e^{(n-1)\phi/2}∂$ [20, §4], [22, §1.4]. The eta invariant is also a conformal invariant [25]. Finally, for $g_{1}$ a smoothing on $N=\partial M\times[0,1]$ of the metric $g|_{\partial M}$ at $t=0$ to a product metric near $t=1,$ we have $\int_{\partial M}P(g)=\int_{N}\hat{A}(\Omega_{g_{1}}).$ We can clearly construct $\tilde{g}_{1}$ conformal to $g_{1}$ on $N$ in the obvious notation. As above, this implies that $\int_{\partial M}P(g)$ is a conformal invariant. Thus $\operatorname{ind}(\not{\partial}_{g})=\operatorname{ind}(\not{\partial}_{\tilde{g}})$.
For the twisted case, we check that
∂̸$\not{\partial}_{\tilde{g}}^{E}(a^{ij}\psi_{j}\otimes s_{i})∂$ $=(e^{-(n-1)\phi/2}\not{\partial}_{g}e^{(n-1)\phi/2}\otimes{\rm Id}+{\rm Id}\otimes\nabla^{E})(a^{ij}\psi_{j}\otimes s_{i})$
$=e^{-(n-1)\phi/2}e_{k}(e^{(n-1)\phi/2}a^{ij})e_{k}\cdot\psi_{j}\otimes s_{i};$
$\left(e^{-(n-1)\phi/2}\not{\partial}_{g}^{E}e^{(n-1)\phi/2}\right)(a^{ij}\psi_{j}\otimes s_{i})$ $=e^{-(n-1)\phi/2}e_{k}(e^{(n-1)\phi/2}a^{ij})e_{k}\cdot\psi_{j}\otimes s_{i}.$
Thus the twisted Dirac operator is also conformally covariant. (This does not use that $\nabla^{E}$ is flat.) This implies that the $\eta$-invariant for ∂̸$\not{\partial}^{E}∂$ is also a conformal invariant. As in the previous paragraph, $\operatorname{ind}(\not{\partial}^{E})$ is therefore a conformal invariant. ∎
Let $(Y,g)$ be a Riemannian spin manifold. The $\xi$-invariant associated to the unitary representation $\alpha$ of $\pi_{1}(Y)$ is
(29) $\xi_{\alpha}(Y,g)=\frac{1}{2}\left({\rm dim}\ \ker(\not{\partial}^{E_{\alpha}}_{g})+\eta^{E_{\alpha}}_{g}(0)\right)-\frac{k}{2}\left({\rm dim}\ \ker(\not{\partial}_{g})+\eta_{g}(0))\right),$
where $\eta_{g}^{E_{\alpha}}(0)$ is the $\eta$-invariant for ∂̸$\not{\partial}_{g}^{E}∂$, and $k$ is the rank of $\alpha.$ For a family of metrics $\{g_{t}\}$, $\xi_{\alpha}(Y,g_{t})\in\mathbb{R}$ is not independent of $t$, but can only jump by an integer amount when the dimension of either ∂̸$\not{\partial}_{g}∂$ or ∂̸$\not{\partial}^{E}_{g}∂$ jumps. Thus the mod $\mathbb{Z}$ reduction $\tilde{\xi}_{\alpha}(Y)$ of $\xi_{\alpha}(Y,g)$ is a smooth invariant of $(Y,\alpha)$. We call $\xi_{\alpha}(Y,g)$ quasi-topological, and now prove that it is an obstruction to $\eta_{1}>0$ on spin manifolds.
Theorem 4.1.
(i) Let $(M^{n},\partial M,g)$ be a spin manifold with boundary with $\eta_{1}>0$, and either totally umbilic boundary or $\zeta_{1}>0$. Assume that $\pi_{1}(\partial M)$ admits a unitary representation $\alpha$ that extends to a unitary representation of $\pi_{1}(M).$ Then $\xi_{\alpha}(\partial M,g)=0.$
(ii) If $(M,g)$ has $R_{g}>0$ and totally umbilic boundary, and we have the representation hypothesis, then $\xi_{\alpha}(\partial M,g)=0.$
Proof.
As usual, (i) implies (ii). For (i), the left hand side of (28) vanishes for $\tilde{g}$ as in Lem. 4.1, and hence for $g$ by Lemma 4.2. $P(g)=P(\Omega_{g},\omega_{g})$ is a an invariant polynomial in the components of the curvature and connection forms of $g$ [15, §3]. On the right hand side, by Corollary 3.4 there is a different metric $g^{\prime}$ conformal to $g$ with ker(∂̸$\not{\partial}_{i^{*}g^{\prime}})=0( ∂$ on $\partial M$. By the conformal invariance of this kernel, ker(∂̸$\not{\partial}_{i^{*}g})=0( ∂$. Therefore,
(30) $\eta_{g}(0)=-2\int_{M}\hat{A}(\Omega_{g})-2\int_{\partial M}P(g)$
is the integral of locally computed curvature and connection terms.
The APS formula for ∂̸$\not{\partial}_{g}^{E}∂$ is
$\operatorname{ind}(\not{\partial}^{E}_{g})$ $=\int_{M}\hat{A}(\Omega_{g})ch(\Omega_{E})+\int_{\partial M}P^{E}(g)-\frac{1}{2}({\rm dim\ ker}(\not{\partial}^{E}_{i^{*}g})+\eta^{E}_{g}(0))$
$=k\int_{M}\hat{A}(\Omega_{g})+k\int_{\partial M}P(g)-\frac{1}{2}({\rm dim\ ker}(\not{\partial}^{E}_{i^{*}g})+\eta^{E}_{g}(0)).$
This follows from
[…]
where $c$ is a dimension constant, in the usual notation for Chern-Simons forms. As in the last paragraph, we can assume that $\operatorname{ind}(\not{\partial}^{E}_{g})=0$. The argument in Corollary 3.4 applies to ∂̸$\not{\partial}^{E}_{i^{*}g}∂$, by the twisted Licherowicz formula (26). Thus we can also assume ${\rm dim\ ker}(\not{\partial}^{E}_{i^{*}g})=0$. We obtain
(31) $\eta^{E}_{g}(0)=-2k\int_{M}\hat{A}(\Omega_{g})-2k\int_{\partial M}P(g).$
From (30), (31), we get
$\xi_{\alpha}(\partial M,g)$ $=\frac{1}{2}\eta^{E_{\alpha}}_{g}(0)-\frac{k}{2}\eta_{g}(0)$
$=-k\left(\int_{M}\hat{A}(\Omega_{g})+\int_{\partial M}P(g)\right)+k\left(\int_{M}\hat{A}(\Omega_{g})+\int_{\partial M}P(g)\right)$
$=0.$
Remark 4.1.
(i) The Theorem implies that the nonvanishing of the topological invariant $\tilde{\xi}_{\alpha}(Y)$ is an obstruction to $\eta_{1}>0$, but in practice computing $\tilde{\xi}_{\alpha}(Y)$ comes down to computing $\xi_{\alpha}(Y,g)$ and reducing mod $\mathbb{Z}.$
(ii) (30) is proved in [4, Thm. 3.9], if $M$ admits a psc metric $g$ which is a product near $\partial M.$ Such a metric has vanishing second fundamental form, so $h_{g}=0$. Thus these metrics have $\eta_{1}>0$, and so form a special case of the Theorem.
5. Manifolds with $3$-dimensional lens space boundaries
There are explicit formulas for the $\tilde{\xi}$-invariants of spherical space forms; these are particularly tractable for lens spaces with odd order fundamental group. Making use of these formulas, we present some concrete examples of spin $4$-manifolds with boundary a disjoint union of lens spaces that cannot admit a metric with $\eta_{1}>0$, and hence have no psc metric with minimal boundary (or a metric with nonnegative scalar curvature and positive mean curvature).
The lens space $L(p,q)$ is the quotient of $S^{3}\subset\mathbb{C}^{2}$ by the action of $\mathbb{Z}_{p}$ generated by
$T\cdot(z_{1},z_{2})=(\omega z_{1},\omega^{q}z_{2}),\ \text{where}\ \omega=e^{2\pi i/p}.$
Here $T$ is a fixed generator of $\mathbb{Z}_{p}$, which can also be viewed as a fixed generator of $\pi_{1}(L)$. The round metric on $S^{3}$ descends to a metric on $L$, which evidently has positive scalar curvature. The $\tilde{\xi}$-invariants with respect to this standard metric have been computed by Gilkey. The precise formula is a little unwieldy, but we only need a non-vanishing result.
Lemma 5.1 (Gilkey).
Let $\alpha:\pi_{1}(L(p,q))\to\operatorname{U}(1)$ take $T$ to $\omega$. Then, with respect to any metric $g$ of positive scalar curvature, $\xi_{\alpha}(L(p,q),g)\neq 0$.
Proof.
For $p$ odd (so there is a unique spin structure), Gilkey [16, Theorem 2.5 (c)] shows that with respect to the standard round metric $g_{0}$ on $L(p,q)$
$\xi_{\alpha}(L(p,q),g_{0})\;=\;-(d/p)\cdot(p+1)/2\pmod{\mathbb{Z}},$
where $d$ is a certain integer relatively prime to $48$ and $p$. One can easily see that the right hand side is never zero modulo the integers, and hence $\xi_{\alpha}(L(p,q),g)$ is not zero as a real number.
A recent result of Bamler and Kleiner [5] shows that the space of psc metrics on any $3$-manifold is contractible, and in particular is path-connected. By the vanishing of the kernel of the Dirac operator on a manifold with a psc metric, there is no spectral flow for the $\eta$-invariant of the Dirac operator along a path of psc metrics. The same holds for the twisted Dirac operator, and so $\xi_{\alpha}$ is constant along a path of psc metrics. It follows that $\xi_{\alpha}(L(p,q),g)$ is non-zero for any psc metric $g$ on $L(p,q)$. ∎
5.1. A rational coboundary for lens spaces
An argument with the Atiyah-Hirzebruch spectral sequence (AHSS) [12] says that the bordism group $\Omega^{\,\operatorname{spin}}_{2k+1}(BG)$ is torsion for any finite group $G$. In other words, for any odd-dimensional spin manifold $Y$ and unitary representation $\alpha$ of $\pi_{1}(Y)$ that factors through a finite group, there is a spin manifold $M$ with a unitary representation $\hat{\alpha}$ of its fundamental group, and an integer $N$ with $\partial(M,\hat{\alpha})=N(Y,\alpha)$.
Here is a brief outline of the argument. The AHSS has $E^{2}_{p,q}\cong H_{p}(BG;\Omega^{\,\operatorname{spin}}_{q})$ and converges to $\Omega^{\,\operatorname{spin}}_{*}(BG)$. It is standard that for a finite group, the rational homology of $BG$ vanishes in degree greater than zero. Moreover, the Anderson-Brown-Peterson calculation of the spin cobordism group [3] (see [28, page 336] for the rational calculation) implies that $\Omega^{\,\operatorname{spin}}_{q}\otimes\mathbb{Q}=0$ when $q$ is odd. It follows that the AHSS for $\Omega^{\,\operatorname{spin}}_{*}\otimes\mathbb{Q}$ collapses immediately at the $E^{2}$ term, and so $\Omega^{\,\operatorname{spin}}_{2k+1}(BG)\otimes\mathbb{Q}=0$.
For the lens spaces with odd order fundamental group, we can give an explicit construction. The multiple $N$ we find is not optimal; for instance, we find a spin $4$-manifold $M$ with $\partial(M,\hat{\alpha})=p^{2}(L(p,q),\alpha)$, whereas a closer examination of the AHSS shows that $\Omega^{\,\operatorname{spin}}_{3}(B\mathbb{Z}_{p})\cong\mathbb{Z}_{p}$ so $N=p$ would do.
The construction begins with the surface $\Sigma_{p}^{0}$ drawn (for $p=3$) in Figure 1, inspired by Figure 1 in [18] and [17, §5]. It may described as a stack of $p$ discs, joined by $p$ half-twisted bands as shown in the figure. The surface $\Sigma_{p}^{0}$ has a $p$-fold symmetry $\tau$, given by counterclockwise rotation by angle $2\pi/p$ with fixed point set $p$ points, one on each of the discs. Furthermore $\partial\Sigma_{p}^{0}$ is $p$ circles, cyclically permuted by $\tau$. Hence the symmetry extends to an action of $\mathbb{Z}_{p}$ on the closed surface
$\Sigma_{p}=\Sigma_{p}^{0}\cup pD^{2}.$
[…]
Figure 1. The surface $\Sigma_{3}^{0}$
Note that the symmetry $\tau$ on $\Sigma_{p}^{0}$ comes from the $p$-fold symmetry of $S^{3}$ given by rotation around the indicated axis. Now $S^{3}$ is a spin manifold, and it is standard that this symmetry lifts to the spin bundle of $S^{3}$. Therefore $\tau$ lifts to the spin bundle of the restriction of the spin structure on $S^{3}$ to $\Sigma_{p}^{0}$.
We claim that this equivariant spin structure $\mathfrak{s}$ extends to $\Sigma_{p}$ as well. To see this, it suffices to show that the induced spin structure on each boundary component of $\Sigma_{p}^{0}$ is the bounding spin structure on $S^{1}$ (and hence extends, uniquely, over each of the $p$ discs that are added.). By construction, $p$ copies of the spin structure on the circle is a boundary, for it bounds $\Sigma_{p}^{0}$. But the spin cobordism group $\Omega_{2}^{\,\operatorname{spin}}$ is $\mathbb{Z}_{2}$, and since $p$ is odd, it follows that $(S^{1},\mathfrak{s})=0$ in $\Omega_{2}^{\,\operatorname{spin}}$. Alternatively, one can note that the normal framing on each boundary component induced by the inward normal of $\Sigma_{p}^{0}$ is $p-1$, which is even. Hence the $4$-manifold obtained by adding $2$-handles to $S^{3}\times I$ along $\partial\Sigma_{p}^{0}$ with framing $p-1$ is spin, and contains $\Sigma_{p}$. The spin structure on this $4$-manifold induces a spin structure on $\Sigma_{p}$ extending the given one on $\Sigma_{p}^{0}$.
To construct a bounding manifold for $p^{2}L(p,q)$, consider the product of two copies of $\Sigma_{p}$. There is an action of $\mathbb{Z}_{p}$ on $(\Sigma_{p})^{2}$ where the generator, say $T$, acts by $\tau$ on the first factor of $\Sigma_{p}$ and by $\tau^{q}$ on the second factor. The fixed point set of $T$ therefore consists of $p^{2}$ isolated points, where the local representation is exactly the action that defines the lens space $L(p,q)$. Now remove an invariant $D^{4}$ neighborhood of each fixed point, and take the quotient by $\mathbb{Z}_{p}$. The quotient is a spin manifold with boundary $p^{2}L(p,q)$. The orientation on the lens space boundaries is opposite to the orientation implicit in the definition, so we set $M=M(p,q)$ to be this spin manifold with orientation reversed. By construction, there is a homomorphism $\hat{\alpha}:\pi_{1}(M)\to\mathbb{Z}_{p}$ that restricts to $\alpha$ on each copy of $L(p,q)$.
Combining Thm. 4.1 with the construction above gives our main example.
Theorem 5.1.
$M(p,q)$ has no metric with $\eta_{1}>0$ and mean convex boundary. In particular, $M(p,q)$ has no psc metric with mean convex (e.g., minimal) boundary.
Proof.
If $M(p,q)$ has a metric $g$ with $\eta_{1,g}>0$ and totally umbilic boundary, then by Cor. 3.3 we can assume that $R_{i^{*}\tilde{g}}>0$ for some $\tilde{g}$ conformal to $g$. By Lem. 5.1, $\xi_{\alpha}(L(p,q),\tilde{g})\neq 0.$ $R_{i^{*}\tilde{g}}>0$ implies $\zeta_{1,\tilde{g}}>0$, so by Thm. 4.1, we have $\eta_{1,\tilde{g}}\leq 0$ and hence $\eta_{1,g}\leq 0$, a contradiction. In particular, $g$ cannot be psc with totally geodesic boundary.
By [7, Thm. 27], a psc metric $g_{1}$ with mean convex boundary can be deformed through psc metrics to a psc metric $g$ with totally geodesic boundary. Since $\eta_{1.g}>0$, $g$ and hence $g_{1}$ cannot exist. ∎
Remark 5.1.
If we replace mean convex boundary with minimal boundary in the theorem, the proof follows from earlier work [1], [2, Thms. 2.1, 4.6], where it is proved that such a metric can be deformed to a product metric near the boundary.
5.2. Higher dimensions
Let $(Y,g,\alpha)$ be a $(4k-1)$-dimensional Riemannian spin manifold with $k>1$, where $\alpha$ is a representation of $\pi_{1}(Y)$ that factors through a finite group. We know $N(Y,\alpha)=\partial(M,\hat{\alpha})$ for some $N$ with $M$ spin and $\hat{\alpha}$ an extension of $\alpha$. If $\xi_{\alpha}(Y,g)\neq 0$, then $M$ admits no metric $h$ with $\eta_{1,h}>0$ and $i^{*}h=g$ on each component of $\partial M.$ Thus the $\xi$-invariant is an obstruction to extending $g$ from $NY$ to a metric $h$ on a bounding manifold $M$ such that $h$ is psc with totally geodesic boundary.
In the quasi-topological direction, we list two cases where $M$ admits no metric with $\eta_{1}>0,$ but we know of no examples.
  1. (1)
    Assume $\tilde{\xi}_{\alpha}(Y)\neq 0\in\mathbb{R}/\mathbb{Z}.$ If a fixed metric $g$ on $Y$ has $\eta_{1,g}>0$, then $\xi_{\alpha}(Y,g)\neq 0\in\mathbb{R}.$ As in Cor. 3.3, we may assume that $i^{*}g$ is psc. If the space of psc metrics is connected, then as in Lem. 5.1, $\xi_{\alpha}(NY,g)=N\xi_{\alpha}(Y,g)\neq 0.$ This is a contradiction. There are examples of higher dimensional lens spaces with nonzero $\xi$-invariants, but we don’t have an example of a higher dimensional manifold with connected space of psc metrics.
  2. (2)
    If the space of psc metrics is not connected, then for $NY=\amalg_{i=1}^{N}Y_{i}$,
    $\xi_{\alpha}(NY,g)=\sum_{i=1}^{N}\xi_{\alpha}(Y_{i},g)$ may vanish, since the $\xi_{\alpha}(Y_{i},g)$ differ by integers. However, this cannot occur if $\xi_{\alpha}(Y_{i},g)\not\in\mathbb{Q}.$ While irrational $\xi$-invariants occur for $Y=S^{1}$, we again know of no higher dimensional examples. In particular, the examples in [8, Lemma 2.3] are all rational.